OpenStudy (anonymous):

use logarithmic differentiation to find the derivative of g(x)= x^(x)^(2).

OpenStudy (anonymous):

implicit differentiation for this one. use logs to get exponents down

OpenStudy (anonymous):

\[x^{x^2}\]?

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

take the log, simplify using properties of the log, then find the derivative, then multiply by the original function

OpenStudy (anonymous):

\[\ln(x^{x^2})=x^2\ln(x)\]

OpenStudy (anonymous):

take the derivative using product rule

OpenStudy (anonymous):

yeup, just what i said :D

OpenStudy (anonymous):

ohhh okay so i dont use the natural log

OpenStudy (anonymous):

the derivative is \[x+2x\ln(x)\] then multiply by original thing and be done

OpenStudy (anonymous):

that is the natural log yes

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw